OpenStudy (anonymous):

Write an equationin standard form for the ellipse with foci (7, 0) and (-7, 0) and y-intercepts 6 and -6. a. x^2/36 + y^2/85 = 7 b. x^2/36 + y^2/85 = 1 c. x^2/85 + y^2/36 = 7 d. x^2/36 + y^2/36 = 1

OpenStudy (mertsj):

\[c ^{2}=a ^{2}-b ^{2}\]

OpenStudy (mertsj):

49=a^2-36 a^2=85 Should be x^2/85+y^2/36=1

OpenStudy (mertsj):

Did you type an answer wrong?

OpenStudy (mertsj):

c=7, b = 6

OpenStudy (anonymous):

no, those answer choices are correct.

OpenStudy (mertsj):

c should say = 1

OpenStudy (mertsj):

Does it say = 1?

OpenStudy (anonymous):

no, c says =7

OpenStudy (mertsj):

Something has to be wrong. The foci are on the x axis which means the major axis is horizontal and (a,0) and (-a,0) are the endpoints of the major axis. The y intercepts are (0,6) and (0,-6) which means that b = 6.

OpenStudy (mertsj):

So a^2 must be 85. The standard form of an ellipse with major axis horizontal and center at origin is x^2/a^2 +y^2/b^2 =1

OpenStudy (mertsj):

So x^2/85 +y^2/36 = 1 is the correct equation.

OpenStudy (anonymous):

Maybe the person that wrote it, wrote it wrong. but I do understand what you mean.